OpenStudy (anonymous):

A train travels 76 kilometers in 4 hours, then travels 63 kilometers in 4 hours. What is its average speed

OpenStudy (aveline):

What is \[\frac{ 76km+63km }{ 2 }\] ?

OpenStudy (anonymous):

\[average ~speed=\frac{ 76+63 }{ 4+4 }=?\]

OpenStudy (aveline):

@surjithayer That's not going to work. You should write it like this: \[\frac{ (76*4)+(63*4) }{ total hours traveled }\]

OpenStudy (anonymous):

if it is written for 4 hours it runs at a speed of 76 kn/ hr and for next 4 hours it runs at a speed of 63 kh/hr. then it can be solved as you have written.